Show that 𝑓(𝑧) = 𝑥 2 + 𝑦 2 where 𝑧 = 𝑥 + 𝑖𝑦 is not analytic anywhere using Cauchy Riemann equations.


1
Expert's answer
2021-11-10T18:07:18-0500

Cauchy Riemann Equations:

Given, "f(x+iy)=u(x,y)+iv(x,y)" and it should satisfy 

"\frac{∂u}

{∂x}=


\frac{∂v}{

∂y}


\\and\\\frac{ ∂u}{

∂y}


=\frac{−∂v}{

∂x}"


If Cauchy Riemann equations are satisfied, then the function is analytic.


"Given\\f(z)=x^2+y^2\\U(x,y)=x^2+y^2=0\\U_x=2x;U_y=2y\\v_x=0;v_y=0"

Cauchy Riemann equations.

"U_x=v_y;U_y=-v_x\\2x=0;2y=0"

Since the given function doesn't satisfy the property of Cauchy Riemann equation which shows

that "f(z)=x^2+y^2\\"

Is not analytic.